Find the area of a parallelogram with side lengths of 8 and 14 feet and an interior angle measure of 40 degrees.
Found 2 solutions by khwang, cleomenius:
Answer by khwang(438)

For a triangle ABC, its area = ab sin C/2
Now two sides 8,14 with angle 40 deg,
the area of the parallelogram = double of the area of the triangle
= 8*14 sin 40 deg = 71.99

Answer by cleomenius(959)

The formula for the area of a parallelogram is b * h
We are given an angle of 40 degrees, so we can construct a right triangle of degrees 40 90 50
To find the height of the parallelogram, you can now use trigonometry;
sin 40 degrees = opposite / hypotenuse

0.6428 = opposite / hypotenuse
= 0.6428 * 8 = ~ 5.14
This is the height.
We can now use the formula b * h. which comes out to 5.14 * 14 = ~ 71.99
Ancient Wizard Cleomenius